OpenStudy (anonymous):

Elijah and Angelina are playing a game that uses two piles of cards. Each of the 12 cards in one pile lists one month of the year. Each of the 7 cards in the other pile lists a day of the week. On each person’s turn, the player randomly chooses one card from each pile. What is the probability of choosing a month beginning with the letter J and a day beginning with the letter T?

OpenStudy (kropot72):

The cards are in two piles. The drawing of a card from one pile has no effect on the drwing of a card from the other pile. The probability of choosing a J and then a T is the product of the separate probabilities:\[\frac{3}{12}\times \frac{2}{7}\] Now you can calculate the answer.